# scale cd/.style=every label/.append style=scale=#1, cells=nodes=scale=#1

### v1

@@ Abstract.
We define the notion of a Brauer pair of a chain complex, extending the notion of a Brauer pair of a $p$-permutation module introduced by Boltje and Perepelitsky. In fact, the Brauer pairs of a splendid Rickard complex $C$ coincide with the set of Brauer pairs of the corresponding $p$-permutation equivalence $\Lambda(C)$ induced by $C$. As a result, we derive structural results for splendid Rickard complexes which are analogous to known structural properties for $p$-permutation equivalences. In particular, we show splendid Rickard equivalences induce local splendid Rickard equivalences between normalizer block algebras as well as centralizer block algebras.
@@
Brauer pairs, splendid Rickard complex, $p$-permutation equivalence, block theory, Broué’s abelian defect group conjecture
@@ 1. Introduction
Broué’s abelian defect group conjecture is one of the hallmark conjectures of modular representation theory. It predicts an “equivalence” between a block with abelian defect group, and a corresponding block associated to the normalizer of the defect group. Moreover, such equivalences are predicted to induce “local equivalences” as well, in that for these block algebras, there are corresponding blocks associated to centralizer subgroups which are also equivalent in the same manner.
@@
Historically, these equivalences were first predicted by Broué to exist as isotypies, a bouquet of character-theoretic equivalences called “perfect isometries” for pairs of local subgroups. However, it was suspected that isotypies were the “shadow” of a categorical equivalence from which all of the components of the isotypy are induced. Rickard refined this conjecture in [16] by introducing “splendid” derived equivalences, objects which are now known as splendid Rickard complexes. Splendid Rickard complexes are chain complexes of $p$-permutation bimodules which have “diagonal” vertices, and induce equivalences between the derived categories corresponding to the block algebras. The key insight of Rickard was that $p$-permutation was the correct property to require for the existence of equivalences between local subgroups. In this case, the Brauer construction, a common construction when working with $p$-permutation modules, induces the local equivalences. Such complexes were later extended to the language of (almost) source algebras by Linckelmann in [11]. In [5], Boltje and Xu introduced $p$-permutation equivalences, a type of equivalence for blocks which exists on the representation ring level. Such equivalences live between isotypies and splendid Rickard complexes, in the sense that in order to induce an isotypy from a splendid Rickard equivalence, one first constructs a corresponding $p$-permutation equivalence by taking the Lefschetz invariant of the chain complex, i.e. taking an alternating sum of terms. Much of the theory of splendid Rickard complexes and $p$-permutation equivalences should be expected to coincide - for instance, to induce local equivalences between centralizer block algebras for both types of equivalences, one applies the Brauer construction in the same manner. However, a splendid Rickard complex has additional structure that a $p$-permutation equivalence does not - namely, the differentials of the chain complex, and a grading of components.
@@
The preliminary notion of $p$-permutation equivalence in [5] made restrictive assumptions, namely that the blocks have a common defect group and that certain fusion systems are equivalent. The main results were obtained by assuming that the defect group was abelian. Boltje and Perepelistky in [4] gave a wider definition for $p$-permutation equivalences, allowing for direct sums of blocks, removing the assumption of a common defect group, and allowing the vertices of any indecomposable summand to be “twisted diagonal.” Under these looser assumptions, it is initially unclear how one would obtain local equivalences, or if the stricter structural properties of a $p$-permutation equivalence would hold in the case of block equivalences. The machinery of Brauer pairs was necessary to resolve these questions, and in fact, many of the structural invariants, such as blocks having isomorphic fusion systems, continue to hold. By applying the Brauer construction to a $p$-permutation equivalence at certain Brauer pairs, one obtains a local $p$-permutation equivalence, leading to the notion of a Brauer pair of a $p$-permutation equivalence. If $A$ and $B$ are direct summands of group algebras $kG$ and $kH$ and $\gamma$ is a $p$-permutation equivalence for $A$ and $B$, we say the pair $(\Delta(P,\phi,Q),e\otimes f^{*})$, where $\Delta(P,\phi,Q)\leq G\times H$ is a twisted diagonal subgroup, $e$ a block of $kC_{G}(P)$, and $f$ a block of $kC_{H}(Q)$, is a $\gamma$-Brauer pair if $e\gamma(\Delta(P,\phi,Q))f\neq 0$.
@@
Boltje and Perepelitisky developed this theory further, deducing the structure of the poset of Brauer pairs for $p$-permutation equivalences. Moreover, they showed that the local equivalences induced extend beyond blocks of centralizer group algebras - they may be extended to stronger equivalences for normalizer blocks as well. A key observation was that applying the Brauer construction to a $p$-permutation equivalence does not immediately lead to a centralizer $p$-permutation equivalence, but rather a module which is “invertible” under an extended tensor product. This suggests that there is more structure which is missed by only considering local centralizer $p$-permutation equivalences.
@@
The goal of this paper is adapting much of the theory in [4] to splendid Rickard complexes. We adopt a similarly flexible definition for splendid Rickard complexes, allowing for twisted diagonal vertices and for the equivalence to be between sums of blocks. We then introduce the notion of a Brauer pair for a splendid Rickard complex. Given a splendid Rickard complex $C$ for direct summands $A$ and $B$ of group algebras $kG$ and $kH$, respectively, we say $(\Delta(P,\phi,Q),e\otimes f^{*})$ is a $C$-Brauer pair if $eC(\Delta(P,\phi,Q))f$ is a noncontractible chain complex. Our main result is the following, which is Corollary 8.8:
@@
Let $C$ be a splendid Rickard complex between $A$ and $B$ and let $\gamma:=\Lambda(C)$ be its corresponding $p$-permutation equivalence. Then $\mathcal{BP}(C)=\mathcal{BP}(\gamma)$. In particular,
@@
We deduce that splendid Rickard complexes induce equivalences at the normalizer level as well. This is Corollary 8.9.
@@
Let $C$ be a splendid Rickard complex between blocks $A\subseteq kG$ and $B\subseteq kH$ and let $(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $C$-Brauer pair. Set $I=N_{G}(P,e)$ and $J=N_{H}(Q,f)$ and let $\hat{e}$ and $\hat{f}$ denote the unique block idempotents of $k[N_{G}(P)]$ and $k[N_{H}(Q)]$ which cover $e$ and $f$, respectively. Then,
@@
is a splendid Rickard complex for blocks algebras $k[N_{G}(P)]\hat{e}$ and $k[N_{H}(Q)]\hat{f}$.
Finally, we deduce that the vertices of the bimodules which appear in an indecomposable splendid Rickard complex $C$ are governed by the Brauer pairs of $C$. This is Theorem 9.5.
@@
Let $C$ be an indecomposable splendid Rickard complex for block algebras $A\subseteq kG$ and $B\subseteq kH$. Every trivial source bimodule occurring in $C$ has a maximal Brauer pair contained in $\mathcal{B}\mathcal{P}(C)$. In particular, the set consisting of all Brauer pairs for every trivial source bimodule appearing in $C$ and the set of $C$-Brauer pairs coincide.
The paper is organized as follows. Section 2 introduces some preliminary definitions involving direct products of groups. Section 3 reviews some prerequisites on $p$-permutation modules and the Brauer construction. Section 4 is a review of block theory, Brauer pairs, and the fusion system of a block. Section 5 establishes the definitions of $p$-permutation equivalences and splendid Rickard complexes we will use for the paper, and proves a few basic results for splendid Rickard complexes which hold for more restrictive definitions. Section 6 reviews preliminaries on extended tensor products with a focus on extended tensor products of $p$-permutation bimodules. Section 7 recalls the definition of a Brauer pair for a $p$-permutation module and a $p$-permutation equivalence as defined by Perepelitsky and Boltje. Section 8 states and proves the main results of the paper regarding the notion of a Brauer pair for a splendid Rickard complex. Finally, Section 9 proves as corollaries a few structural results about the modules which appear in splendid Rickard complexes, which are analogous to results for $p$-permutation equivalences.
Notation and conventions: For this paper, $G$ and $H$ are finite groups, $k$ is an algebraically closed field of prime characteristic $p$, and $\mathcal{O}$ is a complete discrete valuation ring of characteristic 0 with residue field $k$. The symbol $R$ may be used to denote either $\mathcal{O}$ or $k$. All modules are assumed to be finitely generated, and all chain complexes are assumed to be bounded. Given two $RG$-modules $M$ and $N$, the tensor product $M\otimes_{k}N$ is an $RG$-module with $G$-action $g(m\otimes n)=gm\otimes gn$. We denote by $M^{*}:=\mathrm{Hom}_{R}(M,R)$ the $R$-dual of $M$, which is an $RG$-module via $G$-action $g\cdot f(m)=f(g^{-1}m)$. Similarly, if $K$ is a finite group, given a $(RG,RH)$-bimodule $M$ and a $(RH,RK)$-bimodule $N$, the tensor product $M\otimes_{RH}N$ is a $(RG,RK)$-bimodule. We denote by $M^{*}$ the $R$-dual of $M$ which is an $(RH,RG)$-bimodule under the action $h\cdot f(m)\cdot g=f(gmh)$. We write $s_{p}(G)$ to denote the set of $p$-subgroups of $G$. We refer the reader to [13, Sections 1.17] for the constructions of tensor products and internal homs of chain complexes.
@@ 2. Subgroups of direct products of groups
We begin by recalling some constructions related to subgroups of direct products of groups. We refer the reader to [6, Part I, Chapter 2] for more details.
@@
    The subgroup $X^{\circ}:=\{(h,g)\in H\times G\mid(g,h)\in X\}$ is called the opposite subgroup of $X$.
@@
The next lemma is routine to verify.
@@
Let $G$ and $H$ be finite groups, let $\Delta(P,\phi,Q)$ be a twisted diagonal subgroup of $G\times H$, and let $S\leq N_{G}(P)$ and $T\leq N_{H}(Q)$. We denote by $N_{(S,\phi,T)}$ the subgroup of $S$ consisting of all elements $g\in S$ such that there exists an element $h\in T$ satisfying $c_{g}\phi c_{h}=\phi$ as functions from $Q$ to $P$. Note that if $S$ contains $C_{G}(P)$, then also $N_{(S,\phi,T)}$ also contains $C_{G}(P)$. Moreover, if $P\leq S$ and $Q\leq T$, then $P\leq N_{(S,\phi,T)}$. We further set $N_{\phi}:=N_{(N_{G}(P),\phi,N_{H}(Q))}$. Note that $N_{\phi}$ corresponds to $N_{\phi^{-1}}$ in the literature on fusion systems, see for example in [2].
The following are routine to verify, and will be important in the sequel.
@@ 3. $p$-permutation modules and the Brauer construction
Throughout this chapter, let $G$ be a finite group and let $(K,\mathcal{O},k)$ be a $p$-modular system large enough for $G$. That is $\mathcal{O}$ and $k$ have $|G|_{p^{\prime}}$ roots of unity. Unless otherwise specified, $R$ will denote either $\mathcal{O}$ or $k$. The properties listed in this section are well-known, we refer the reader to [13] or [8] for more details and proofs.
@@
      $M$ is isomorphic to a direct sum of permutation modules.
@@
    The functor induced by the surjection $\mathcal{O}\to k$, ()¯$\overline{(-)}:{}_{\mathcal{O}G}\mathbf{mod}\to{}_{kG}\mathbf{mod}$ induces a vertex-preserving bijection from the set of isomorphism classes of $p$-permutation $\mathcal{O}G$-modules to the set of isomorphism classes of $p$-permutation $kG$-modules.
@@
For this reason, indecomposable $p$-permutation modules are called trivial source modules. We denote the full subcategory of $p$-permutation $RG$-modules by ${}_{RG}\textbf{triv}$, and the full subcategory of $p$-permutation $(RG,RH)$-bimodules by ${}_{RG}\textbf{triv}_{RH}$. Given an idempotent $e\in Z(RG)$, we may consider ${}_{RGe}\textbf{triv}$ a full subcategory of ${}_{RG}\textbf{triv}$.
@@
This may be realized as a functor $(-)(P):{}_{kG}\textbf{mod}\to{}_{k[N_{G}(P)/P]}\textbf{mod}$ as well, via
@@
The following diagram of functors commutes for any $p$-subgroup $P\in s_{p}(G)$:
{tikzcd}
If $P$ is not a $p$-group, then $M(P)=0$ for every $RG$-module $M$. For this reason, we restrict our attention to taking Brauer quotients at $p$-groups. The Brauer construction restricts to a functor $(-)(P):{}_{RG}\mathbf{triv}\to{}_{k[N_{G}(P)/P]}\mathbf{triv}$. The Brauer construction is an additive functor, but in general it is not exact. Since it is additive, it can be extended to a functor of chain complexes, $\operatorname{Ch}({}_{RG}\textbf{mod})\to\operatorname{Ch}({}_{k[N_{G}(P)/P]}\textbf{mod})$ and similarly for ${}_{RG}\mathbf{triv}$.
@@
    Let $M$ be a $RG$-module, $P\in s_{P}(G)$, and $x\in G$. Then there is a natural isomorphism (${}^{x}(M(P))\cong M({}^{x}P)$ of $k[{}^{x}(N_{G}(P)/P)]$-modules.
The Brauer construction behaves especially well with respect to $p$-permutation modules. The following properties are well-known, and are collected from [13, §5.10-5.11].
@@
    Let $M,N$ be $p$-permutation $RG$-modules and $P\in s_{p}(G)$. Then there is a natural isomorphism $(M\otimes_{R}N)(P)\cong M(P)\otimes_{k}N(P)$ of $k[N_{G}(P)/P]$-modules.
@@
    Let $Q\trianglelefteq P\in s_{p}(G)$ and let $M$ be a $p$-permutation $RG$-module. Then $M(P)\cong M(Q)(P)$, regarding $M(Q)$ as a $kN_{G}(Q)$-module. In particular, if $M(P)\neq 0$, then $M(R)\neq 0$ for any subgroup $R\leq P$.
@@
    For each $P\in s_{p}(G)$, the Brauer construction $M\mapsto M(P)$ induces a bijection between the set of isomorphism classes of indecomposable $p$-permutation $RG$-modules with vertex $P$ and the set of isomorphism classes of indecomposable projective $k[N_{G}(P)/P]$-modules. Moreover, if $M$ is indecomposable with vertex $P$, then $M(P)$ is its Green correspondent.
@@
$\operatorname{br}_{P}=\operatorname{br}^{G}_{P}:RG\to kC_{G}(P),\quad\sum_{g\in G}\alpha_{g}\cdot g\mapsto\sum_{g\in C_{G}(P)}\overline{\alpha_{g}}\cdot g.$
@@
We refer the reader to [2, Part IV] for more details on Brauer maps. The final proposition allows us to compute the Brauer construction of a module that has been cut by a block idempotent.
@@
[4, Lemma 3.8] Let $M\in{}_{RG}\mathbf{triv}$, $P\in s_{p}(G)$, and $i\in(kG)^{H}$ an idempotent, fixed under the conjugation action of $H\leq G$ with $P\leq H$. Then for each $m\in M^{P}$, one has $\operatorname{Br}^{M}_{P}(im)=\operatorname{br}^{G}_{P}(i)\operatorname{Br}^{M}_{P}(m)$. In particular, one obtains a natural isomorphism $(iM)(P)\xrightarrow{\sim}\operatorname{br}^{G}_{P}(i)M(P)$ of $kN_{H}(P)$-modules.

### v2

# Brauer pairs for splendid Rickard equivalences

@@ Abstract.
We define the notion of a Brauer pair of a chain complex, extending the notion of a Brauer pair of a $p$-permutation module introduced by Boltje and Perepelitsky. In fact, the Brauer pairs of a splendid Rickard equivalence $C$ coincide with the set of Brauer pairs of the corresponding $p$-permutation equivalence $\Lambda(C)$ induced by $C$. As a result, we derive structural results for splendid Rickard equivalences that correspond to known structural properties for $p$-permutation equivalences. In particular, we show splendid Rickard equivalences induce local splendid Rickard equivalences between normalizer block algebras as well as centralizer block algebras.
@@
Brauer pair, splendid Rickard equivalence, $p$-permutation equivalence, block theory, abelian defect group conjecture
@@ 1. Introduction
Broué’s abelian defect group conjecture, one of the hallmark conjectures of modular representation theory, predicts an equivalence between a block with abelian defect group and a corresponding “local” block associated to the normalizer of the defect group. Moreover, these equivalences should induce “local equivalences;” for these block algebras, there are corresponding pairs of blocks associated to centralizers which are also equivalent.
@@
Historically, these equivalences were first predicted by Broué to exist either as isotypies, a bouquet of character-theoretic equivalences called perfect isometries for pairs of local subgroups, or as derived equivalences, which then induced a perfect isometry. However, Broué suspected that isotypies were the shadow of a categorical equivalence from which all of the components of the isotypy are induced. Rickard refined this conjecture in [Ri96] by introducing “splendid” derived equivalences, which are now often called splendid Rickard equivalences. These equivalences are chain complexes of $p$-permutation bimodules which have “diagonal” vertices, and induce not only derived equivalences, but equivalences between homotopy categories of $p$-permutation modules for the block algebras as well. In fact, recent work of Balmer and Gallauer in [BG23a] demonstrates that these homotopy categories of $p$-permutation modules admit the corresponding derived module categories as Verdier quotients. Rickard’s key insight was that $p$-permutation is the correct property to require for the existence of equivalences between local subgroups. In this case, the Brauer construction, a common construction when working with $p$-permutation modules, induces the local equivalences. Such complexes were later extended to the language of (almost) source algebras by Linckelmann in [Li98].
In [BoXu07], Boltje and Xu introduced $p$-permutation equivalences, a type of equivalence for blocks which exists on the representation ring level. Such equivalences live between isotypies and splendid Rickard equivalences, in the sense that in order to induce an isotypy from a splendid Rickard equivalence, one first constructs a corresponding $p$-permutation equivalence by taking the Lefschetz invariant of the chain complex, i.e. taking an alternating sum of terms. Much of the theory of splendid Rickard equivalences and $p$-permutation equivalences should be expected to coincide - for instance, to induce local equivalences between centralizer block algebras for both types of equivalences, one applies the Brauer construction in the same manner. However, a splendid Rickard equivalence has additional structure that a $p$-permutation equivalence does not, those of differentials and grading.
@@
The earliest notions of $p$-permutation equivalence in [BoXu07] made restrictive assumptions, namely that the blocks have a common defect group and that certain block fusion systems are equivalent. Boltje and Perepelistky in [BoPe20] gave a wider definition for $p$-permutation equivalences, allowing for direct sums of blocks, removing the assumption of a common defect group, and allowing the vertices of any indecomposable summand to be “twisted diagonal.” Under these looser assumptions, it is initially unclear how one would obtain local equivalences, or if the stricter structural properties of a $p$-permutation equivalence would hold in the case of block equivalences. The machinery of Brauer pairs was necessary to resolve these questions, and in fact, many of the structural invariants, such as blocks having isomorphic fusion systems, continue to hold. By applying the Brauer construction to a $p$-permutation equivalence at certain Brauer pairs, one obtains a local $p$-permutation equivalence, leading to the notion of a Brauer pair of a $p$-permutation equivalence. If $A$ and $B$ are direct summands of group algebras $kG$ and $kH$ and $\gamma$ is a $p$-permutation equivalence for $A$ and $B$, we say the pair $(\Delta(P,\phi,Q),e\otimes f^{*})$, where $\Delta(P,\phi,Q)\leq G\times H$ is a twisted diagonal subgroup, $e$ a block of $kC_{G}(P)$, and $f$ a block of $kC_{H}(Q)$, is a $\gamma$-Brauer pair if $e\gamma(\Delta(P,\phi,Q))f\neq 0$.
@@
Boltje and Perepelitsky developed this theory further, deducing the structure of the poset of Brauer pairs for $p$-permutation equivalences. Moreover, they showed that the local equivalences induced extend beyond blocks of centralizer group algebras - they may be extended to stronger equivalences for normalizer blocks as well. A key observation was that applying the Brauer construction to a $p$-permutation equivalence does not immediately lead to a centralizer $p$-permutation equivalence, but rather a module which is “invertible” under an extended tensor product. This suggests that there is more structure which is missed by only considering local centralizer $p$-permutation equivalences.
@@
The goal of this paper is adapting much of the theory in [BoPe20] to splendid Rickard equivalences. We adopt a similarly flexible definition for splendid Rickard equivalences, allowing for twisted diagonal vertices and for the equivalence to be between sums of blocks. We then introduce the notion of a Brauer pair for a splendid Rickard equivalence. Given a splendid Rickard equivalence $C$ for direct summands $A$ and $B$ of group algebras $kG$ and $kH$, respectively, we say $(\Delta(P,\phi,Q),e\otimes f^{*})$ is a $C$-Brauer pair if $eC(\Delta(P,\phi,Q))f$ is a noncontractible chain complex. Our main result is the following, which is Corollary 8.8:
@@
Let $C$ be a splendid Rickard equivalence between $A$ and $B$ and let $\gamma:=\Lambda(C)$ be its corresponding $p$-permutation equivalence. Then $\mathcal{BP}(C)=\mathcal{BP}(\gamma)$. In particular,
@@
These results were deduced by Boltje and Perepelitsky in [BoPe20] in the context of Brauer pairs for $p$-permutation equivalences; our main contribution is showing that the notion of Brauer pairs agree. Part (d) recovers a classical result of Puig; see for instance [Pu99]. We deduce that splendid Rickard equivalences induce equivalences at the normalizer level as well. This is Corollary 8.9.
@@
Let $C$ be a splendid Rickard equivalence for blocks $A\subseteq kG$ and $B\subseteq kH$ and let $(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $C$-Brauer pair. Set $I=N_{G}(P,e)$ and $J=N_{H}(Q,f)$ and let $\hat{e}$ and $\hat{f}$ denote the unique block idempotents of $k[N_{G}(P)]$ and $k[N_{H}(Q)]$ which cover $e$ and $f$, respectively. Then,
@@
is a splendid Rickard equivalence for block algebras $k[N_{G}(P)]\hat{e}$ and $k[N_{H}(Q)]\hat{f}$.
Finally, we deduce that the vertices of the bimodules which appear in an indecomposable splendid Rickard equivalence $C$ are governed by any single maximal Brauer pair for $C$. This is Corollary 9.5.
@@
Let $C$ be an indecomposable splendid Rickard equivalence for block algebras $A\subseteq kG$ and $B\subseteq kH$ and let $(\Delta(D,\phi,E),e\otimes f^{*})$ be a maximal $C$-Brauer pair. Every trivial source bimodule occurring in $C$ has a maximal Brauer pair contained in $(\Delta(D,\phi,E),e\otimes f^{*})$. In particular, the set consisting of all Brauer pairs for every trivial source bimodule appearing in $C$ and the set of $C$-Brauer pairs coincide.
The paper is organized as follows. Section 2 introduces some preliminary definitions involving direct products of groups. Section 3 reviews some prerequisites on $p$-permutation modules and the Brauer construction. Section 4 is a review of block theory, Brauer pairs, and the fusion system of a block. Section 5 establishes the definitions of $p$-permutation equivalences and splendid Rickard equivalences we will use for the paper, and proves a few basic results for splendid Rickard equivalences which hold for more restrictive definitions. Section 6 reviews preliminaries on extended tensor products with a focus on extended tensor products of $p$-permutation bimodules. Section 7 recalls the definition of a Brauer pair for a $p$-permutation module and a $p$-permutation equivalence as defined by Perepelitsky and Boltje. Section 8 states and proves the main results of the paper regarding the notion of a Brauer pair for a splendid Rickard equivalence. Finally, Section 9 proves as corollaries a few structural results about the modules which appear in splendid Rickard equivalences, which are analogous to results for $p$-permutation equivalences.
Notation and conventions: For this paper, $G$ and $H$ are finite groups, $k$ is an algebraically closed field of prime characteristic $p$, and $\mathcal{O}$ is a complete discrete valuation ring of characteristic 0 with residue field $k$. The symbol $R$ may be used to denote either $\mathcal{O}$ or $k$. All modules are assumed to be finitely generated, and all chain complexes are assumed to be bounded. Given two $RG$-modules $M$ and $N$, the tensor product $M\otimes_{k}N$ is an $RG$-module with $G$-action $g(m\otimes n)=gm\otimes gn$. We denote by $M^{*}:=\mathrm{Hom}_{R}(M,R)$ the $R$-dual of $M$, which is an $RG$-module via $G$-action $g\cdot f(m)=f(g^{-1}m)$. Similarly, if $K$ is a finite group, given a $(RG,RH)$-bimodule $M$ and a $(RH,RK)$-bimodule NN, the tensor product $M\otimes_{RH}N$ is a $(RG,RK)$-bimodule. We denote by $M^{*}$ the $R$-dual of $M$ which is an $(RH,RG)$-bimodule under the action $h\cdot f(m)\cdot g=f(gmh)$. We write $s_{p}(G)$ to denote the set of $p$-subgroups of $G$. We refer the reader to [Li181, Sections 1.17] for the constructions of tensor products and internal homs of chain complexes.
@@ 2. Subgroups of direct products of groups
We begin by recalling some constructions related to subgroups of direct products of groups. We refer the reader to [Bo10, Part I, Chapter 2] for more details.
@@
    The subgroup $X^{\circ}:=\{(h,g)\in H\times G\mid(g,h)\in X\}$ is called the subgroup opposite to $X$.
@@
The following lemma consists of routine verifications.
@@
Let $G$ and $H$ be finite groups, let $\Delta(P,\phi,Q)$ be a twisted diagonal subgroup of $G\times H$, and let $S\leq N_{G}(P)$ and $T\leq N_{H}(Q)$. We denote by $N_{(S,\phi,T)}$ the subgroup of $S$ consisting of all elements $g\in S$ such that there exists an element $h\in T$ satisfying $c_{g}\phi c_{h}=\phi$ as functions from $Q$ to $P$. Note that if $S$ contains $C_{G}(P)$, then also $N_{(S,\phi,T)}$ also contains $C_{G}(P)$. Moreover, if $P\leq S$ and $Q\leq T$, then $P\leq N_{(S,\phi,T)}$. We further set $N_{\phi}:=N_{(N_{G}(P),\phi,N_{H}(Q))}$. Note that $N_{\phi}$ corresponds to $N_{\phi^{-1}}$ in the literature on fusion systems, see for example [AKO11].
The following lemma is again routine.
@@ 3. $p$-permutation modules and the Brauer construction
Throughout this chapter, let $G$ be a finite group and let $(K,\mathcal{O},k)$ be a $p$-modular system with $\mathcal{O}$ and $k$ having $|G|_{p^{\prime}}$ roots of unity. Unless otherwise specified, $R$ will denote either $\mathcal{O}$ or $k$. The properties listed in this section are well-known, we refer the reader to [Br85] or [Li181] for more details and proofs.
@@
      Let $S\leq G$ be a Sylow $p$-subgroup of $G$. Then $\operatorname{Res}^{G}_{S}M$ is a permutation module.
@@
    The functor induced by the surjection $\mathcal{O}\to k$, $\overline{(-)}:{}_{\mathcal{O}G}\mathbf{mod}\to{}_{kG}\mathbf{mod}$ induces a vertex-preserving bijection from the set of isomorphism classes of $p$-permutation $\mathcal{O}G$-modules to the set of isomorphism classes of $p$-permutation kGkG-modules.
@@
For this reason, indecomposable $p$-permutation modules are called trivial source modules. We denote the full subcategory of $p$-permutation RGRG-modules by $_{RG}\textbf{triv}$, and the full subcategory of $p$-permutation $(RG,RH)$-bimodules by $_{RG}\textbf{triv}_{RH}$. Given an idempotent $e\in Z(RG)$, we may consider $_{RGe}\textbf{triv}$ a full subcategory of $_{RG}\textbf{triv}$.
@@
The Brauer construction is also considered as a functor $(-)(P):{}_{kG}\textbf{mod}\to{}_{k[N_{G}(P)/P]}\textbf{mod}$ as well, via
@@
This choice is convention is minor, as the following diagram of functors commutes for any $p$-subgroup $P\in s_{p}(G)$:
$_{\mathcal{O}G}\textbf{mod}$$_{kG}\textbf{mod}$$_{kN_{G}(P)/P}\textbf{mod}$$\scriptstyle{\overline{(-)}}$$\scriptstyle{-(P)}$$\scriptstyle{-(P)}$
[…]
@@
[…]
The Brauer construction behaves especially well with respect to $p$-permutation modules. The following properties are well-known, and are collected from [Li181, §5.10-5.11].
@@
    Let $M,N$ be $p$-permutation RGRG-modules and $P\in s_{p}(G)$. Then there is a natural isomorphism $(M\otimes_{R}N)(P)\cong M(P)\otimes_{k}N(P)$ of $k[N_{G}(P)/P]$-modules.
@@
    Let $Q\trianglelefteq P\in s_{p}(G)$ and let $M$ be a $p$-permutation RGRG-module. Then $M(P)\cong M(Q)(P)$, regarding $M(Q)$ as a $kN_{G}(Q)$-module. In particular, if $M(P)\neq 0$, then $M(R)\neq 0$ for any subgroup $R\leq P$.
@@
    For each $P\in s_{p}(G)$, the Brauer construction $M\mapsto M(P)$ induces a bijection between the set of isomorphism classes of indecomposable $p$-permutation RGRG-modules with vertex $P$ and the set of isomorphism classes of indecomposable projective $k[N_{G}(P)/P]$-modules. Moreover, if $M$ is indecomposable with vertex $P$, then $M(P)$ is its Green correspondent.
@@
$\operatorname{br}_{P}=\operatorname{br}^{G}_{P}:(RG)^{P}\to kC_{G}(P),\quad\sum_{g\in G}\alpha_{g}\cdot g\mapsto\sum_{g\in C_{G}(P)}\overline{\alpha_{g}}\cdot g.$
@@
We refer the reader to [AKO11, Part IV] for more details on Brauer maps. The final proposition allows us to compute the Brauer construction of a module that has been cut by a block idempotent.
@@
[BoPe20, Lemma 3.8] Let $M\in{}_{RG}\mathbf{triv}$, $P\in s_{p}(G)$, and $i\in(kG)^{H}$ an idempotent, fixed under the conjugation action of $H\leq G$ with $P\leq H$. Then for each $m\in M^{P}$, one has $\operatorname{Br}^{M}_{P}(im)=\operatorname{br}^{G}_{P}(i)\operatorname{Br}^{M}_{P}(m)$. In particular, one obtains a natural isomorphism $(iM)(P)\xrightarrow{\sim}\operatorname{br}^{G}_{P}(i)M(P)$ of $kN_{H}(P)$-modules.